OpenStudy (anonymous):

evaluate if a=2 and b=-5 (a + b)^2

OpenStudy (anonymous):

9

OpenStudy (anonymous):

9

OpenStudy (angela210793):

[2+(-5)]^2=(2-5)^2=(-3)^2=9